OpenStudy (anonymous):

If 75 g of FeCO3 is heated with excess Oxygen, 45 g of Fe2O3 is produced. what is the percent yeild?????

OpenStudy (unklerhaukus):

ok first lets balance the chemical equation

OpenStudy (unklerhaukus):

we only need to worry about this sub reaction \[\text{FeCO}_3\to \text{Fe}_2\text O_3\]

OpenStudy (anonymous):

its 4FeCO3 + O2 --->2Fe2O3 + 4CO2 thats what the book gives me

OpenStudy (unklerhaukus):

\[ 4\text{FeCO}_3\to 2\text{Fe}_2\text O_3\] ok so the thing to notice is ratio of two \[\color{brown} 2\text{FeCO}_3\to \text{Fe}_2\text O_3\]

OpenStudy (unklerhaukus):

Two moles of reactant will theoretically form one mole of Product

OpenStudy (unklerhaukus):

Can you find how many moles of reactant 75 g of FeCO_3 is ?

OpenStudy (unklerhaukus):

\[\boxed{ n=\dfrac mM}\] \(n\) is number of moles \([\text{moles}]\) \(m\) is mass \([\text g]\) \(M\) is molar mass \(\frac{[\text g]}{[\text{moles}]}\)

OpenStudy (unklerhaukus):

\[M_{\text{FeCO}_3}=M_{\text{Fe}}+M_{\text{C}}+3M_{\text{O}}\]

OpenStudy (unklerhaukus):

\[M_{\text{Fe}}=55.85\\M_{\text{C}}=12.01\\M_{\text{O}}=16.00\]

OpenStudy (unklerhaukus):

are you following? @needsomuchhelp

OpenStudy (unklerhaukus):

the relative percentage yield is \[\frac{\text{Actual yield}}{\text{Theoretic yield}}\times100\%\]